OpenStudy (haley.fieldss):

Can some one simply 5x+2(x+6)?

likeabossssssss (likeabossssssss):

what d u think it is ?

OpenStudy (mhchen):

what do you think it is?

KjSaif (kjsaif):

^^^^^

OpenStudy (maddielb):

PEMDAS, distribute 2(x+6) and get 5x+2x+12, combine like terms, 7x+12

OpenStudy (haley.fieldss):

thanks. I was right

KjSaif (kjsaif):

5x+2(x+6) Distribute: =5x+(2)(x)+(2)(6) =5x+2x+12 Combine Like Terms: =5x+2x+12 =(5x+2x)+(12) =7x+12

OpenStudy (maddielb):

You're welcome

likeabossssssss (likeabossssssss):

do u got any more?

KjSaif (kjsaif):

^^

OpenStudy (haley.fieldss):

Yeah

likeabossssssss (likeabossssssss):

post them in a new post tho thank u

KjSaif (kjsaif):

^^